OpenStudy (billy-bob):

I need help to find a solution set for y = -1/3x +2

OpenStudy (maxwellfish):

What does the question specifically say?

OpenStudy (billy-bob):

Explain how you determined where to draw the line and shade the area that represents the solution set.

OpenStudy (maxwellfish):

The line fo the slope or for what

OpenStudy (billy-bob):

the line of the slope

OpenStudy (maxwellfish):

m= -1/3

OpenStudy (billy-bob):

Thank you!

OpenStudy (maxwellfish):

No Problem

OpenStudy (maxwellfish):

Not trying to be rude but could you medal my response please

OpenStudy (billy-bob):

Oh yes sorry i forgot lol

OpenStudy (maxwellfish):

Thank You

OpenStudy (billy-bob):

No problem!

OpenStudy (princeharryyy):



OpenStudy (princeharryyy):

@Billy-Bob

OpenStudy (billy-bob):

i thought it was y=mx+b

OpenStudy (princeharryyy):

it doesn't matter u write that a, b or c. they are all constants

OpenStudy (billy-bob):

true

OpenStudy (princeharryyy):

:)

OpenStudy (billy-bob):

:)

OpenStudy (billy-bob):

thank you!

OpenStudy (princeharryyy):

yw :)

OpenStudy (mathmale):

The equation y = -1/3x +2 should, I suspect, be written as y = -(1/3)x +2. This is much less ambiguous than is y = -1/3x +2 . This function has an infinite number of solutions. YOU choose an x value and calculate y from the function. Then write your result in the form (x,y). Doing this will give you ONE point on the line represented by y = -(1/3)x +2 . You could find a few more points that lie on the line, and then quit.

OpenStudy (billy-bob):

Thank you!